A Gaussian surface in the shape of a right circular cylinder with end caps has a radius of 17.1 cm and a length of 77.9 cm. Through one end there is an inward magnetic flux of 25.4 μWb. At the other end there is a uniform magnetic field of 2.40 mT, normal to the surface and directed outward. What are the (a) magnitude and (b) direction (inward or outward) of the net magnetic flux through the curved surface?
Answer #1

Part a The total magnetic flux is: 0 = 01 +02 In this case: 01 = -25.4uWb (inward) 02 = +AB (outward) $2 = TR2B = 1(0.171)?(2